# On the strength of connectedness of unions of random graphs

### v1

@@ 2.3 Proof of Theorem 2
Now we evaluate ${\bf{E}}N_{k}$ and ${\bf{E}}N_{*k}$. To this aim use approximations (43) and (44) shown in Lemma 5 below. Fix $v\in{\cal V}$. We have, by symmetry,
@@ 3.2 Inequalities related to expansion property
${\bf{P}}\{{\cal A}_{1,k,\ell}\}={\bf{P}}\{1\in V_{k,\ell}^{*}\}={\bf{P}}\{v^{*}\in V_{k,\ell}^{*}\}={\bf{E}}({\bf{P}}\{v^{*}\in V_{k,\ell}^{*}|F_{k,\ell}^{*}\})=\frac{v_{k,\ell}}{n}.$
In the last step we evaluated the conditional probability ${\bf{P}}\{v^{*}\in V_{k,\ell}^{*}|F_{k,\ell}^{*}\}=\frac{v_{k,\ell}}{n}$.
@@
$1-{\bf{P}}\{{\cal A}_{2,k,\ell}\}$ $={\bf{P}}\{\{1,2\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}={\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}$
$={\bf{E}}\left({\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}|F_{k,\ell}^{*}\}\right)=\frac{k}{\binom{n}{2}}$
In the last step we used the fact that ${\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}|F_{k,\ell}^{*}\}=\frac{k}{\binom{n}{2}}$. Furthermore, the inequality $2k\leq v_{k,\ell}$ implies
${\bf{P}}\{{\cal A}_{2,k,\ell}\}=1-\frac{k}{\binom{n}{2}}\geq 1-\frac{v_{k,\ell}}{(n)_{2}}.$
Finally, a calculation shows that for $2\leq v_{k,\ell}\leq n$ the right side is lower bounded by $v_{k,\ell}\frac{2}{n}\left(1-\frac{2}{n}\right)-\frac{1}{2}v^{2}_{k,\ell}\frac{2^{2}}{n^{2}}.$
@@
We write event ${\cal A}_{r,k,\ell}$ in the form $\left(\cup_{i}{\cal L}_{i}\right)\cup\left(\cup_{j}{\cal T}_{j}\right)$ and apply the inclusion-exclusion inequalities
$S_{L}+S_{T}-Q\leq{\bf{P}}\{{\cal A}_{r,k,\ell}\}={\bf{P}}\left\{\left(\bigcup_{i}{\cal L}_{i}\right)\bigcup\left(\bigcup_{j}{\cal T}_{j}\right)\right\}\leq S_{L}+S_{T},$ (61)
@@
Proof of (62), (63). We firstly bound probabilities ${\bf{P}}\{{\cal L}_{1}\}$ and ${\bf{P}}\{{\cal T}_{1}\}$ from below
${\bf{P}}\{{\cal L}_{1}\}$ $=\frac{r(n-r)}{\binom{n}{2}}=2\frac{r(n-r)}{n(n-1)}>2\frac{r(n-r)}{n^{2}}=2\frac{r}{n}-2\frac{r^{2}}{n^{2}},$
${\bf{P}}\{{\cal T}_{1}\}=1-{\bf{P}}\{{\bar{\cal T}}_{1}\}$ $=1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{r^{3}}{n^{3}}-\frac{(n-r)^{3}}{n^{3}}=3\frac{r}{n}-3\frac{r^{2}}{n^{2}}.$
Next, we bound ${\bf{P}}\{{\cal L}_{1}\}$ and ${\bf{P}}\{{\cal T}_{1}\}$ from above
${\bf{P}}\{{\cal L}_{1}\}$ $=2\frac{r(n-r)}{n(n-1)}\leq 2\frac{r}{n},$
${\bf{P}}\{{\cal T}_{1}\}$ $<1-\frac{(n-r)_{3}}{(n)_{3}}=\frac{(n)_{3}-(n-r)_{3}}{(n)_{3}}$
$=r\frac{(n)_{2}+n(n-2-r)+(n-1-r)_{2}}{(n)_{3}}$
$\leq 3\frac{r}{n}.$
Proof of (64), (65), (66). To show inequality (64) we apply the product rule
${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}$ $={\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}=\frac{r(n-r)}{\binom{n}{2}}\frac{r(n-r)-1}{\binom{n}{2}-1}<\left(\frac{r(n-r)}{\binom{n}{2}}\right)^{2}$
$=({\bf{P}}\{{\cal L}_{1}\})^{2}=\left(\left(2\frac{r}{n}-2\frac{r^{2}}{n^{2}}\right)\left(1+\frac{1}{n-1}\right)\right)^{2}.$
The proof of (65), (66) is a bit more involved. Let $\mathbb{L}$} denote the set edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$. Let $\mathbb{T}$} denote the set of open triangles $\tau$ of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$. Given a set $A$ we denote by ${\cal K}_{A}$ the complete graph on the vertex set $A$.
Proof of (65) We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$. Furthermore, given $T_{1}^{*}$, the random edge $L_{1}^{*}$ is uniformly distributed across the set of edges of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$. We partition ${\mathbb{T}}={\mathbb{T}}_{1}\cup{\mathbb{T}}_{2}\cup{\mathbb{T}}_{3}\cup{\mathbb{T}}_{4}$ so that
@@
${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{r}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
To prove (65) we show that for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$
${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal L}_{1}\}.$ (68)
Proof of (68). We have ${\bf{P}}\{{\cal L}_{1}\}=\frac{N_{L}}{M_{L}}$, where $N_{L}=|{\mathbb{L}}|=(n-r)r$ stands for the number of edges in the set $\mathbb{L}$ and $M_{L}=\binom{n}{2}$ is the number of edges of ${\cal K}_{n}$. For $\tau\in{\mathbb{T}}_{i}$ we have ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$, where $N_{i}$ is the number of edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$ that are non-incident to $\tau$; $M_{i}$ is the number of edges of ${\cal K}_{n}$ that are non-incident to $\tau$. To prove (68) we show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{L}}{M_{L}}$ for $i=1,2,3,4$. Note that $M_{i}=\binom{n-3}{2}$ for each $i=1,2,3,4$.
Case $i=1$. We have $N_{1}=(n-r-1)(r-2)$. Inequalities
@@
imply that for $r\leq n/3$ we have
$\frac{N_{1}}{M_{1}}=\frac{N_{L}}{M_{L}}\frac{N_{1}}{N_{L}}\frac{M_{L}}{M_{1}}<\frac{N_{L}}{M_{L}}.$
Case $i=2$. We have $N_{2}=(n-r-2)(r-1)$. Inequalities
$\frac{N_{2}}{N_{L}}$ $=\frac{(n-r-2)(r-1)}{(n-r)r}=\left(1-\frac{2}{n-r}\right)\left(1-\frac{1}{r}\right)<1-\frac{1}{r},$
$\frac{M_{2}}{M_{L}}$ $>1-\frac{6}{n}$
imply that for $r\leq n/6$ we have
$\frac{N_{2}}{M_{2}}=\frac{N_{L}}{M_{L}}\frac{N_{2}}{N_{L}}\frac{M_{L}}{M_{2}}<\frac{N_{L}}{M_{L}}.$
The remaining cases $i=3,4$ are treated in much the same way, since $N_{3}=N_{2}$, $N_{4}=N_{1}$.
Proof of (66). We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ and, given $T_{1}^{*}$, the random open triangle $T_{2}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$. We have, by the total probability formula,
${\bf{P}}\{{\cal T}_{2}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{r}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
To prove (66) we show that for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$
${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal T}_{2}\}.$ (69)
Proof of (69). We have ${\bf{P}}\{{\cal T}_{2}\}=\frac{N_{T}}{M_{T}}$, where $N_{T}=|{\mathbb{T}}|=3\binom{r}{2}(n-r)+3\binom{n-r}{2}r$ is the number of triangles in the set $\mathbb{T}$ and $M=3\binom{n}{3}$ is the number of open triangles $\nu$ of ${\cal K}_{n}$. For $\tau\in{\mathbb{T}}_{i}$ we have ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$, where $N_{i}$ is the number of open triangles $\nu$ of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$ that are non-incident to $\tau$; $M_{i}$ denote the number of open triangles $\nu$ of ${\cal K}_{n}$ that are non-incident to $\tau$. To prove (69) we show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{T}}{M_{T}}$ for $i=1,2,3,4$. Note that $M_{i}=3\binom{n-3}{3}$ for each $i=1,2,3,4$.
Case $i=1$. We have
$N_{1}=3\binom{r-2}{2}(n-r-1)+3\binom{n-r-1}{2}(r-2)=3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta,$
where
@@
$N_{1}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)=\max\{\alpha,\beta\}N_{T}<\left(1-\frac{2}{r}\right)N_{T}.$
Furthermore, we have
$M_{1}$ $=\frac{(n-3)_{3}}{(n)_{3}}M_{T}=\left(1-\frac{3}{n}\right)\left(1-\frac{3}{n-1}\right)\left(1-\frac{3}{n-2}\right)M_{T}$ (70)
$>\left(1-\frac{3}{n}-\frac{3}{n-1}-\frac{3}{n-2}\right)M_{T}>\left(1-\frac{10}{n}\right)M_{T}.$
The last inequality holds for $n\geq 11$. Combining these inequalities we upperbound the ratio
$\frac{N_{1}}{M_{1}}<\frac{1-2r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}.$
For $r\leq n/5$ we obtain $\frac{N_{1}}{M_{1}}<\frac{N_{T}}{M_{T}}$.
Case $i=2$. We have
$N_{2}=3\binom{r-1}{2}(n-r-2)+3\binom{n-r-2}{2}(r-1)=3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta,$
where
@@
$N_{2}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)=\max\{\alpha,\beta\}N_{T}<\left(1-\frac{1}{r}\right)N_{T}.$
Recall that $M_{2}=M_{1}$. Invoking (70) we obtain
$\frac{N_{2}}{M_{2}}<\frac{1-r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}.$
For $r\leq n/10$ we obtain $\frac{N_{2}}{M_{2}}<\frac{N_{T}}{M_{T}}$.
The remaining cases $i=3,4$ are treated in much the same way, since $N_{3}=N_{2}$, $N_{4}=N_{1}$. ∎

### v2

@@ 1 Introduction and results
Condition (5) can be relaxed in the sense that the strict identity ${\bf{E}}X_{n,i_{*}}(t)=0$ on the right of (5) can be replaced by a weaker condition that $\max_{1\leq t<a}{\bf{E}}X_{n,i_{*}}(t)$ tends to 0 sufficiently fast as $n\to+\infty$. It seems that the rate of $\ln^{-1}n$ would suffice.
@@ 2.3 Proof of Theorem 2
Now we evaluate ${\bf{E}}N_{k}$ and ${\bf{E}}N_{*k}$. To this aim we use approximations (43) and (44) shown in Lemma 5 below. Fix $v\in{\cal V}$. We have, by symmetry,
@@ 3.2 Inequalities related to expansion property
${\bf{P}}\{{\cal A}_{1,k,\ell}\}={\bf{P}}\{1\in V_{k,\ell}^{*}\}={\bf{P}}\{v^{*}\in V_{k,\ell}^{*}\}=\frac{v_{k,\ell}}{n}.$
@@
$1-{\bf{P}}\{{\cal A}_{2,k,\ell}\}$ $={\bf{P}}\{\{u^{*},v^{*}\}\in\{L^{*}_{1},\dots,L^{*}_{k}\}\}+{\bf{P}}\{\{u^{*},v^{*}\}\cap V_{k,\ell}^{*}=\emptyset\}$
$=\frac{k}{\binom{n}{2}}+\frac{(n-v_{k,\ell})_{2}}{(n)_{2}}.$
Now, using $2k\leq v_{k,\ell}$ one easily shows (60).
@@
We write event ${\cal A}_{r,k,\ell}$ in the form ${\cal A}_{r,k,\ell}=\left(\cup_{i}{\cal L}_{i}\right)\cup\left(\cup_{j}{\cal T}_{j}\right)$ and apply the inclusion-exclusion inequalities
$S_{L}+S_{T}-Q\leq{\bf{P}}\{{\cal A}_{r,k,\ell}\}\leq S_{L}+S_{T},$ (61)
@@
Proof of (62). We evaluate ${\bf{P}}\{{\cal L}_{1}\}=2\frac{r(n-r)}{n(n-1)}$ and invoke inequalities
$2\frac{r}{n}\geq 2\frac{r(n-r)}{n(n-1)}>2\frac{r(n-r)}{n^{2}}=2\frac{r}{n}-2\frac{r^{2}}{n^{2}},$
Proof of (63). We evaluate ${\bf{P}}\{{\cal T}_{1}\}=1-{\bf{P}}\{{\bar{\cal T}}_{1}\}=1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}$ and invoke inequalities
$3\frac{r}{n}\geq 1-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{(r)_{3}}{(n)_{3}}-\frac{(n-r)_{3}}{(n)_{3}}>1-\frac{r^{3}}{n^{3}}-\frac{(n-r)^{3}}{n^{3}}=3\frac{r}{n}-3\frac{r^{2}}{n^{2}}.$
Proof of (64). We apply the product rule ${\bf{P}}\{{\cal L}_{1}\cap{\cal L}_{2}\}={\bf{P}}\{{\cal L}_{1}\}{\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}$ and estimate
${\bf{P}}\{{\cal L}_{2}|{\cal L}_{1}\}=\frac{(r-1)(n-r-1)}{\binom{n-2}{2}}<\frac{r(n-r)}{\binom{n}{2}}={\bf{P}}\{{\cal L}_{1}\}.$
The inequality above is equivalent to the inequality
$\left(1-\frac{1}{r}\right)\left(1-\frac{1}{n-r}\right)<\left(1-\frac{2}{n}\right)\left(1-\frac{2}{n-1}\right),$
which follows from the inequality $1-\frac{1}{r}<\left(1-\frac{2}{n}\right)\left(1-\frac{2}{n-1}\right)$ valid for $r\leq n/4$.
Proof of (65), (66). Let $\mathbb{L}$} denote the set edges of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$. Let $\mathbb{T}$} denote the set of open triangles $\tau$ of ${\cal K}_{n}$ connecting $[r]$ and $n]\setminus[r]$. Given a set $A$ we denote by ${\cal K}_{A}$ the complete graph on the vertex set $A$.
Proof of (65). We recall that $T_{1}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$. Furthermore, given $T_{1}^{*}$, the random edge $L_{1}^{*}$ is uniformly distributed across the set of edges of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$. We partition ${\mathbb{T}}={\mathbb{T}}_{1}\cup{\mathbb{T}}_{2}\cup{\mathbb{T}}_{3}\cup{\mathbb{T}}_{4}$ so that
@@
${\bf{P}}\{{\cal L}_{1}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{4}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
To prove (65) we show that ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal L}_{1}\}$ for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$.
Denote, for short, $N_{L}=|{\mathbb{L}}|=r(n-r)$ and $M_{L}=\binom{n}{2}$ so that ${\bf{P}}\{{\cal L}_{1}\}=\frac{N_{L}}{M_{L}}$. Similarly, for $\tau\in{\mathbb{T}}_{i}$ we denote by $M_{i}$ (and $N_{i}$) the number of edges of ${\cal K}_{n}$ nonincident to $\tau$ (and connecting $[r]$ and $n]\setminus[r]$) so that ${\bf{P}}\{{\cal L}_{1}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$. Note that $M_{i}=\binom{n-3}{2}$ for $i=1,2,3,4$. We show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{L}}{M_{L}}$ for $i=1,2,3,4$ and $2\leq r\leq n/6$.
Let $i=1$. We have $N_{1}=(n-r-1)(r-2)$. Inequalities
@@
imply $\frac{N_{1}}{M_{1}}<\frac{N_{L}}{M_{L}}$. Let $i=2$. We have $N_{2}=(n-r-2)(r-1)$. Inequalities
$\frac{N_{2}}{N_{L}}=\frac{(n-r-2)(r-1)}{(n-r)r}=\left(1-\frac{2}{n-r}\right)\left(1-\frac{1}{r}\right)<1-\frac{1}{r}$
and $\frac{M_{2}}{M_{L}}=\frac{M_{1}}{M_{L}}>1-\frac{6}{n}$ shown above imply $\frac{N_{2}}{M_{2}}<\frac{N_{L}}{M_{L}}$. The remaining cases $i=3,4$ are treated in much the same way.
Proof of (66). We have, by the total probability formula,
${\bf{P}}\{{\cal T}_{2}\cap{\cal T}_{1}\}$ $=\sum_{\tau\in{\mathbb{T}}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}=\sum_{i=1}^{4}\sum_{\tau\in{\mathbb{T}}_{i}}{\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}{\bf{P}}\{T^{*}_{1}=\tau\}.$
To prove (66) we show that ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}\leq{\bf{P}}\{{\cal T}_{2}\}$ for each $i=1,2,3,4$ and every $\tau\in{\mathbb{T}}_{i}$. In the proof we use the fact that given $T_{1}^{*}$, the random open triangle $T_{2}^{*}$ is uniformly distributed across the set of open triangles of ${\cal K}_{n}$ that are non-incident to $T_{1}^{*}$.
Let $N_{T}=|{\mathbb{T}}|=3\binom{r}{2}(n-r)+3\binom{n-r}{2}r$ denote the number of triangles in $\mathbb{T}$ and $M_{T}=3\binom{n}{3}$ denote the number of open triangles of ${\cal K}_{n}$. We have ${\bf{P}}\{{\cal T}_{2}\}=\frac{N_{T}}{M_{T}}$. Similarly, for $\tau\in{\mathbb{T}}_{i}$ we denote by $M_{i}$ (and $N_{i}$) the number of open triangles of ${\cal K}_{n}$ non-incident to $\tau$ (and connecting $[r]$ and $n]\setminus[r]$) so that ${\bf{P}}\{{\cal T}_{2}|T^{*}_{1}=\tau\}=\frac{N_{i}}{M_{i}}$. Note that $M_{i}=3\binom{n-3}{3}$ for each $i=1,2,3,4$. We show that $\frac{N_{i}}{M_{i}}\leq\frac{N_{T}}{M_{T}}$ for $i=1,2,3,4$.
Let $i=1$. We evaluate the number $N_{1}=3\binom{r-2}{2}(n-r-1)+3\binom{n-r-1}{2}(r-2)$ and write it in the form $3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta$, where
@@
$N_{1}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)<\left(1-\frac{2}{r}\right)N_{T}.$
Combining this inequality with the inequality
$M_{1}=\frac{(n-3)_{3}}{(n)_{3}}M_{T}>\left(1-\frac{10}{n}\right)M_{T},$ (68)
which holds for $n\geq 11$, we obtain for $n\geq 11$ and $r\leq n/5$
$\frac{N_{1}}{M_{1}}<\frac{1-2r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}\leq\frac{N_{T}}{M_{T}}.$
Let $i=2$. We evaluate $N_{2}=3\binom{r-1}{2}(n-r-2)+3\binom{n-r-2}{2}(r-1)$ and write it in the form $3\binom{r}{2}(n-r)\alpha+3\binom{n-r}{2}r\beta$, where
@@
$N_{2}\leq\max\{\alpha,\beta\}\left(3\binom{r}{2}(n-r)+3\binom{n-r}{2}r\right)<\left(1-\frac{1}{r}\right)N_{T}.$
This inequality combined with inequality $\frac{M_{2}}{M_{T}}=\frac{M_{1}}{M_{T}}>1-\frac{10}{n}$, see (68) above, implies for $n\geq 11$ and $r\leq n/10$
$\frac{N_{2}}{M_{2}}<\frac{1-r^{-1}}{1-10n^{-1}}\frac{N_{T}}{M_{T}}\leq\frac{N_{T}}{M_{T}}.$
The remaining cases $i=3,4$ are treated in much the same way. ∎